# Multiobjective Aerodynamic Design Optimization of the NASA Common Research Model

### v1

@@ 4.2 The NASA CRM experiment
The NASA common research model (CRM) [53] is an aircraft geometry made for industry-relevant CFD research. We consider a “clean wing” (that is, without nacelles and pylons) version of the CRM geometry, as shown in Figure 5. We leverage the SU2 [18] open-source multiphysics code to perform an inviscid Euler simulation of the flow past the CRM. The computational mesh contains $6.12$M volume elements. A scalar upwind scheme was used for the numerical method with the Venkatakrishnan limiter [54]. A constant CFL number of $10$ was used after it was deemed adequate for satisfactory convergence. Our full SU2 input files and mesh are publicly available at https://github.com/csdlpsu/su2scrips/tree/main/parallel_test.
We consider a two-objective $K=2$ optimization problem, identifying two cruise points as summarized in Table 1; $C_{D1}$ and $C_{D2}$ denote the corresponding drag coefficients at the two cruise points. The optimization problem we will be solving is
$\begin{split}\min_{\mathbf{x}\in\mathcal{X}}&\quad\{C_{D1}(\textbf{x}),C_{D2}(\textbf{x})\},\\ \end{split}$ (8)
@@
We parameterize the wing with a free-form deformation (FFD) [46] box with 162 control points, $24$ of which are allowed to deform – these are the design variables in our optimization problem; Figure 5 displays the FFD box, highlighting the design variables. For control points nearer to the center of the wing, the deformation bounds were (0, 1e-2), slightly farther from the center were (0, 1e-3), and control points near the root and the tip had bounds of (0, 1e-5) to prevent CFD solver divergence.
The CFD simulations were run on Penn State’s Roar Collab supercomputer. We use 6 compute nodes with $48$ cores each for a total of $288$ cores and 2400 GB of memory. For GP training, we first provide $200$ seed points. We run the optimization for an additional $200$ iterations with $q=1$. We compare $q\texttt{POTS}$  to qNEHVI, qPAREGO, which were the best-performing approaches from synthetic experiments, and Sobol random sampling using the hypervolume metric. Due to the computational cost of the CRM experiment, the experiments are not repeated and our comparison is limited to only qNEHVI and qPAREGO.
We begin by presenting the hypervolume history and the Pareto frontiers identified by $q\texttt{POTS}$, in Figure 6. The left panel shows a scatterplot of the objectives chosen by $q\texttt{POTS}$ at each iteration. This reveals an important hypothesis made at the beginning of this manuscript – that the objectives are correlated in parts of the design space while being conflicting in the other parts. Crucially, the conflicting part is where the Pareto optimal designs are also found, emphasizing the importance of multiobjective optimization. The right panel of the figure shows the hypervolume histories of $q\texttt{POTS}$ alongside competing methods – notably, $q\texttt{POTS}$ achieves the highest hypervolume and is the fastest to converge.
@@
The scatterplots of objective values are further shown for the other methods compared against, in Figure 7. Notice that, similar to $q\texttt{POTS}$, all methods identify distinct regions of correlated and conflicting objectives. Further, EHVI and Sobol identify only two objective values that are Pareto optimal. While ParEGO, similar to $q\texttt{POTS}$, identifies several Pareto optimal objectives, the Pareto optimal objectives from ParEGO are dominated by those identified by $q\texttt{POTS}$. In other words, $q\texttt{POTS}$ is able to find designs with higher hypervolume compared to other ParEGO (and the rest of the methods) – an observation that is also confirmed from the right panel of Figure 6.
@@
Next, we compare the pressure coefficient distribution on the wings and the wing cross-sectional shapes identified by $q\texttt{POTS}$. We compare them against an undeformed baseline wing to emphasize the impact of optimization; note that this is the unmodified CRM wing without any optimization involved. Figure 8 shows the pressure coefficients of the undeformed baseline wing on the left and an optimized wing at a Pareto optimal point on the right, for three different Pareto optimal designs; note that these correspond to cruise point 1. The drag coefficients at the two cruise points are printed below the wing as $[C_{D,1},C_{D,2}]$. First, notice that there is a strong shock toward the trailing edge of the wing for the baseline design. A common thread across all three Pareto optimal designs is that the optimized wing, for cruise point $1$, has a lower drag by $45-50$ counts. On the other hand, the optimized drag at cruise point 2 is practically the same (or marginally higher) than the baseline drag for cruise point 2. This shows that the designs at the two cruise points are conflicting – in other words, an independent optimization at one of the design points is likely to result in an undesirably large drag for the other cruise point, thereby emphasizing the need for multiobjective optimization.
(a) Pareto optimal design $1$.
(b) Pareto optimal design $2$.
(c) Pareto optimal design $3$.
@@
We next compare the pressure coefficient at selected spanwise cross-sections in Figure 9; note that this figure shows only the Pareto optimal shapes predicted by $q\texttt{POTS}$ since we already established that $q\texttt{POTS}$ outperforms other methods. In this plot, the red and blue lines correspond to the optimized and baseline wings, respectively. The relocation of the shock (from the trailing edge to leading edge of the wing) and its reduction in strength is quite evident from these plots. Furthermore, one can appreciate the overall reduction in the wave drag component by visualizing the area under these pressure coefficient plots. Crucially, unlike what could have happened in single objective optimization, multiobjective optimization does not entirely eliminate the shock. This is due to the tradeoffs that must be respected at the other operating condition which, thus, again emphasizes the need for multiobjective optimization for aerodynamic design.
@@
Finally, we show the cross-sectional shapes for the Pareto optimal design versus the baseline design in Figure 10. Notice that at $5$ and $95$ wing span, the change in shape due to optimization is subtle, but the impact on the pressure distribution (Figure 9) is a lot more noticeable – this demonstrates the sensitivity of the pressure coefficient to wing shape at the root and tip of the wing. In the remaining locations, the wing shape changes more noticeably, which includes a change in the overall wing twist. In summary, multiobjective optimization identifies a design that offers a pragmatic compromise between the design performance at the two cruise points. Solving this problem efficiently and accurately was enabled by our proposed method $q\texttt{POTS}$.
@@ 5 Conclusion
Aircraft aerodynamic design optimization is better served by designing for the varying operating conditions at multiple sub-segments along the cruise segment, as opposed to considering a fixed operating condition. While existing work addresses this limitation by performing a multi-point optimization that considers a weighted combination of the objectives at all the sub-segments, such an approach depends on specifying appropriate weights and can miss out on identifying other, potentially more promising, designs. In contrast, this work establishes the necessity of solving a multiobjective optimization problem that simultaneously considers the objectives corresponding to all operating conditions. We showed that this is necessary to properly address the conflicts that typically exist between designs optimized individually for each sub-segment.
@@
However, existing methods are prohibitive when applied to multiobjective optimization problems with expensive high-fidelity models. In this regard, we propose a novel multiobjective Bayesian optimization method, called $q\texttt{POTS}$, that is inspired by Thompson sampling, involving Gaussian process surrogates. Specifically, our method adaptively selects a sequence of iterates that are Pareto optimal according to a posterior Gaussian process surrogate model. This way, our method hybridizes classical evolutionary algorithms—which are accurate but not sample efficient—with state-of-the-art Bayesian optimization—which is sample efficient but not accurate enough or scalable—to result in a more accurate and sample efficient approach. We demonstrate our approach on several synthetic experiments as well as the aerodynamic design optimization of the NASA common research model. In all the experiments, we show that our proposed approach either performs the best or is amongst the best performing.
@@
As a first step, this work considered the NASA CRM experiment with two objectives ($K=2$) and without constraints. Therefore, the first leg of future work will consider the inclusion of constraints and more sub-segments considered (higher $K$). We show in the Appendix that extensions to include constraints are straightforward in our approach. Another line of future work is to benchmark our method on the NASA CRM with batch acquisition $(q>1)$ to enable asynchronous parallel evaluations on high-performance computing environments. This is necessary to minimize the wallclock times in the overall optimization involving expensive objectives and constraints. Finally, we also believe in exploring the utility of nonstationary surrogate models, e.g., deep Gaussian process models [40, 41, 4, 5], within our framework. Our proposed framework is generic and replacing the surrogate model with another probabilistic surrogate model is straightforward.

### v2

@@ 4.2 The NASA CRM experiment
The NASA common research model (CRM) [vassberg2008development] is an aircraft geometry made for industry-relevant CFD research. We consider a “clean wing” (that is, without nacelles and pylons) version of the CRM geometry, as shown in Figure˜6. We leverage the SU2 [economon2016su2] open-source multiphysics code to perform an inviscid Euler simulation of the flow past the CRM. The computational mesh contains 6.126.12M volume elements. A scalar upwind scheme was used for the numerical method with the Venkatakrishnan limiter [venkatakrishnan1995convergence]. A constant CFL number of $10$ was used after it was deemed adequate for satisfactory convergence. Our full SU2 input files and mesh are publicly available at https://github.com/csdlpsu/su2scrips/tree/main/parallel_test.
We consider a two-objective $K=2$ optimization problem, identifying two cruise points as summarized in Table˜2; $C_{D1}$ and $C_{D2}$ denote the corresponding drag coefficients at the two cruise points. The optimization problem we will be solving is
[…]
@@
We parameterize the wing with a free-form deformation (FFD) [sederberg1986free] box with 162 control points, $24$ of which are allowed to deform – these are the design variables in our optimization problem; Figure˜6 displays the FFD box, highlighting the design variables. For control points nearer to the center of the wing, the deformation bounds were (0, 1e-2), slightly farther from the center were (0, 1e-3), and control points near the root and the tip had bounds of (0, 1e-5) to prevent CFD solver divergence.
The CFD simulations were run on Penn State’s Roar Collab supercomputer. We use 6 compute nodes with $48$ cores each for a total of $288$ cores and 2400 GB of memory. For GP training, we first provide $200$ seed points. We run the optimization for an additional $200$ iterations with $q=1$. We compare $q\texttt{POTS}$  to qNEHVI, qPAREGO, which were the best-performing approaches from synthetic experiments, and Sobol random sampling using the hypervolume metric. Due to the computational cost of the CRM experiment the experiments are not repeated, and our comparison is limited to only qNEHVI and qPAREGO. Note that BO is particularly suited for optimization of expensive functions where the goal is sample efficiency; therefore, replicating the real-world experiments is not practical. On the other hand, the synthetic experiments are repeated multiple times to establish statistical significance of the reported results. This is standard practice in the Bayesian optimization community [Liang2021, Shields2021, morita2022applying]. Finally, we don’t turn on the Nyström approximation for the CRM experiment. This is because the acquisition computations were still substatially faster than qNEHVI and qPAREGO in this high-dimensional case.
We begin by presenting the hypervolume history and the Pareto frontiers identified by $q\texttt{POTS}$, in Figure˜7. The left panel shows a scatterplot of the objectives chosen by $q\texttt{POTS}$ at each iteration. This reveals an important hypothesis made at the beginning of this manuscript – that the objectives are nonconflicting in parts of the design space while being conflicting in the other parts. Crucially, the conflicting part is where the Pareto optimal designs are also found, emphasizing the importance of multiobjective optimization. The right panel of the figure shows the hypervolume histories of $q\texttt{POTS}$ alongside competing methods – notably, $q\texttt{POTS}$ achieves the highest hypervolume and is the fastest to converge.
@@
The scatterplots of objective values are further shown for the other methods compared against, in Figure˜8. Notice that, similar to $q\texttt{POTS}$, all methods identify distinct regions of nonconflicting and conflicting objectives. Further, EHVI and Sobol identify only two objective values that are Pareto optimal. While ParEGO, similar to $q\texttt{POTS}$, identifies several Pareto optimal objectives, the Pareto optimal objectives from ParEGO are dominated by those identified by $q\texttt{POTS}$. In other words, $q\texttt{POTS}$ is able to find designs with higher hypervolume compared to other ParEGO (and the rest of the methods) – an observation that is also confirmed from the right panel of Figure˜7. The aerodynamic interpretation of the Pareto optimal designs are better obtained from the pressure distribution and wing shape plots discussed next.
@@
(a) Pareto optimal design 11.
(b) Pareto optimal design 22.
(c) Pareto optimal design 33.
@@
Next, we compare the pressure coefficient distribution on the wings and the wing cross-sectional shapes identified by $q\texttt{POTS}$. We compare them against an undeformed baseline wing to emphasize the impact of optimization; note that this is the unmodified CRM wing without any optimization involved. Figure˜9 shows the pressure coefficients of the undeformed baseline wing on the left and an optimized wing at a Pareto optimal point on the right, for three different Pareto optimal designs; note that these correspond to cruise point 1. The drag coefficients at the two cruise points are printed below the wing as $C_{D,1},C_{D,2}]$. First, notice that there is a strong shock toward the trailing edge of the wing for the baseline design. A common thread across all three Pareto optimal designs is that the optimized wing, for cruise point $1$, has a lower drag by 455045-50 counts. On the other hand, the optimized drag at cruise point 2 is practically the same (or marginally higher) than the baseline drag for cruise point 2. This shows that the designs at the two cruise points are conflicting – in other words, an independent optimization at one of the design points is likely to result in an undesirably large drag for the other cruise point, thereby emphasizing the need for multiobjective optimization.
@@
We next compare the pressure coefficient at selected spanwise cross-sections in Figure˜10; note that this figure shows only the Pareto optimal shapes predicted by $q\texttt{POTS}$ since we already established that $q\texttt{POTS}$ outperforms other methods. In this plot, the red and blue lines correspond to the optimized and baseline wings, respectively. The relocation of the shock (from the trailing edge to leading edge of the wing) and its reduction in strength is quite evident from these plots. Furthermore, one can appreciate the overall reduction in the wave drag component by visualizing the area under these pressure coefficient plots. Crucially, unlike what could have happened in single objective optimization, multiobjective optimization does not entirely eliminate the shock. This is due to the tradeoffs that must be respected at the other operating condition which, thus, again emphasizes the need for multiobjective optimization for aerodynamic design.
@@
Finally, we show the cross-sectional shapes for the Pareto optimal design versus the baseline design in Figure˜11. Notice that at $5$ and $95$ % wing span, the change in shape due to optimization is subtle, but the impact on the pressure distribution (Figure˜10) is a lot more noticeable – this demonstrates the sensitivity of the pressure coefficient to wing shape at the root and tip of the wing. In the remaining locations, the wing shape changes more noticeably, which includes a change in the overall wing twist. In summary, multiobjective optimization identifies a design that offers a pragmatic compromise between the design performance at the two cruise points. Solving this problem efficiently and accurately was enabled by our proposed method $q\texttt{POTS}$.
@@ 5 Conclusion
Aircraft aerodynamic design optimization is better served by designing for the varying operating conditions at multiple sub-segments along the cruise segment, as opposed to considering a fixed operating condition. While existing work addresses this limitation by performing a multi-point optimization that considers a weighted combination of the objectives at all the sub-segments, such an approach depends on specifying appropriate weights and can miss out on identifying other, potentially more promising designs. In contrast, this work establishes the necessity of solving a multiobjective optimization problem that simultaneously considers the objectives corresponding to all operating conditions. We showed that this is necessary to properly address the conflicts that typically exist between designs optimized individually for each sub-segment.
@@
However, existing methods are prohibitive when applied to multiobjective optimization problems with expensive high-fidelity models. To address this, we propose a novel multiobjective Bayesian optimization method, called $q\texttt{POTS}$, that is inspired by Thompson sampling, involving Gaussian process surrogates. Specifically, our method adaptively selects a sequence of iterates that are Pareto optimal according to a posterior Gaussian process surrogate model. This way, our method hybridizes classical evolutionary algorithms—which are accurate but not sample efficient—with state-of-the-art Bayesian optimization—which is sample efficient but not accurate enough or scalable—to result in a more accurate and sample efficient approach. We demonstrate our approach on several synthetic experiments as well as the aerodynamic design optimization of the NASA common research model. In all the experiments, we show that our proposed approach either performs the best or is among the best performing.
Known limitations. Despite the strengths of $q\texttt{POTS}$, we acknowledge a few limitations. First, the there is no automatic way to handle high-dimensional scaling which is important for aerodynamic design. However, high-dimensional scaling is a focused research direction in Bayesian optimization [rana2017high, zhang2019high, malu2021bayesian]. However, $q\texttt{POTS}$ is compatible with approaches that exist to address this, e.g., trust-region based multiobjective Bayesian optimization [daulton2022multi] – this is a direction that we plan on pursuing in the future. Second, the posterior sampling complexity is well addressed via the Nyström approximation; however, a well-defined strategy to control the cost-accuracy tradeoff in the algorithm performance doesn’t yet exist. We plan on addressing this as well in the future.
@@
As a first step, this work considered the NASA CRM experiment with two objectives ($K=2$) and without constraints. Therefore, the first leg of future work will consider the inclusion of constraints and more sub-segments considered (higher $K$). We show in the Appendix that extensions to include constraints are straightforward in our approach. Another line of future work is to benchmark our method on the NASA CRM with batch acquisition $(q>1)$ to enable asynchronous parallel evaluations on high-performance computing environments. This is necessary to minimize the wallclock times in the overall optimization involving expensive objectives and constraints. Finally, we also believe in exploring the utility of nonstationary surrogate models, e.g., deep Gaussian process models [rajaram2020deep, rajaram2021empirical, booth2024actively, booth2025contour], within our framework. Our proposed framework is generic and replacing the surrogate model with another probabilistic surrogate model is straightforward.
@@ Acknowledgements
Appendix
5.1 Extension to constrained problems
Our method seamlessly extends to constrained multiobjective problems, including both inequality and equality constraints, and the key step is explained as follows. We demonstrate this with the setting that we have a total of $C$ constraints $\{c_{1},\ldots,c_{C}\}$, and $\mathcal{I}\cup\mathcal{E}=[C]$. Further, let $Y_{k+i\in\mathcal{I}}$ and $Y_{k+i\in\mathcal{E}}$ denote the posterior GP of the $i$th inequality constraint and equality constraint functions, respectively. This way, we fit a total of $K+C$ posterior GPs $Y_{1},\ldots,Y_{K},Y_{K+1},\ldots,Y_{K+C}$. For all $i\in[C]$, we define the constrained $q\texttt{POTS}$  to choose points according to
$\begin{split}p_{\mathcal{X}^{*}}(\mathbf{x})=\int\delta&\left(\mathbf{x}-\operatorname*{argmax}_{\mathbf{x}\in\mathcal{X}}\left[\{Y_{1}(\mathbf{x}),\ldots,Y_{K}(\mathbf{x})\}\times\mathds{1}_{\{\bigcap_{i}Y_{k+i\in\mathcal{E}}(\mathbf{x})=0\}}\times\mathds{1}_{\{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})\geq 0\}}\right]\right)\\ &p(Y_{1}|\boldsymbol{\mathcal{D}}_{n})\ldots p(Y_{K}|\boldsymbol{\mathcal{D}}_{n})dY_{1}\ldots dY_{K},\end{split}$ (9)
where $\mathds{1}_{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})\geq 0}$ and $\mathds{1}_{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})=0}$ are the indicator functions for joint feasibility with inequality and equality constraints, respectively, that take a value of $1$ when feasible and 0 otherwise. This is equivalent to drawing samples from the GP posteriors, filtering them out by the indicator functions, and finding a Pareto optimal solution set per the posterior GP sample paths.
$\begin{split}X^{*}=\operatorname*{argmax}_{\mathbf{x}\in\mathcal{X}}&\left[\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\}\times\mathds{1}_{\{\bigcap_{i}Y_{k+i\in\mathcal{E}}(\mathbf{x})=0\}}\times\mathds{1}_{\{\bigcap_{i}Y_{k+i\in\mathcal{I}}(\mathbf{x})\geq 0\}}\right],\end{split}$ (10)
The rest of the $q\texttt{POTS}$ algorithm remains exactly the same.
5.2 Proof that the multipoint optimum is a member of the Pareto frontier
Let $\mathcal{X}$ be the domain of $f$ and $\boldsymbol{f}:\mathcal{X}\to\mathbb{R}^{m}$ be the vector of objectives to be minimized. Denote by
$F\;=\;\{\,f(\mathbf{x}):\mathbf{x}\in\mathcal{X}\,\}\subset\mathbb{R}^{m}$
the attainable objective set. Let the definition for Pareto optimality presented in Section˜2 hold. Then, we proceed in two steps. First, we show that the weighted-sum minimizers (multi-point optima) are Pareto optimal. Then, we show that the converse is not necessarily always true.
Proposition 5.1.
Let the weights in the multi-point optimum $w\in\mathbb{R}^{m}$ satisfy $w_{i}>0$ for every $i=1,\dots,m$. If
$\mathbf{x}^{\star}\in\arg\min_{\mathbf{x}\in\mathcal{X}}w^{\top}f(\mathbf{x}),$
then $\mathbf{x}^{\star}$ is Pareto optimal.
Proof.
Assume, toward contradiction, that $\mathbf{x}^{\star}$ is not Pareto optimal. Then there exists $y\in\mathcal{X}$ with $f_{i}(\mathbf{y})\leq f_{i}(\mathbf{x}^{\star})$ for all $i$ and $f_{j}(\mathbf{y})<f_{j}(\mathbf{x}^{\star})$ for some index $j$. Multiplying componentwise by the strictly positive weights $w_{i}$ and summing yields
$w^{\top}f(\mathbf{y})\;=\;\sum_{i=1}^{m}w_{i}f_{i}(\mathbf{y})\;<\;\sum_{i=1}^{m}w_{i}f_{i}(\mathbf{x}^{\star})\;=\;w^{\top}f(\mathbf{x}^{\star}),$
which contradicts the minimality of $x^{\star}$. Hence no such $\mathbf{y}$ exists and $\mathbf{x}^{\star}$ must be Pareto optimal. ∎
Converse: when does every Pareto point arise from a weighted sum?
The converse statement ("for every Pareto point there exists $w>0$ such that the point minimizes $w^{\top}f$") is not true in general. A classical sufficient condition that guarantees the converse is convexity of the attainable objective set.
Proposition 5.2 (Converse under convexity).
If the attainable objective set $F$ is convex, then for every Pareto optimal point $f^{\star}\in F$ there exists $w\in\mathbb{R}^{m}$ with $w_{i}\geq 0$, $w\neq 0$, such that $f^{\star}$ minimizes $w^{\top}y$ over $y\in F$. If additionally $f^{\star}$ is a strict Pareto boundary point and appropriate regularity holds, one may choose $w_{i}>0$.
Sketch.
For a convex set $F\subset\mathbb{R}^{m}$ and a boundary point $f^{\star}\in\partial F$, the supporting hyperplane theorem guarantees the existence of a nonzero vector $w$ and scalar $\alpha$ with
$w^{\top}y\geq\alpha=w^{\top}f^{\star}\quad\forall y\in F.$
Thus $f^{\star}$ minimizes $w^{\top}y$ over $F$. Standard arguments translate this supporting hyperplane into a positive-weighted minimization if the Pareto point is not singular with respect to any objective (see multiobjective optimization texts for a full statement). ∎
Consequently, linear (weighted) scalarization is sufficient but not necessary to generate Pareto points unless $F$ is convex. Nonconvex portions of the Pareto front can therefore be missed by any linear scalarization.
5.3 Impact of observation noise variance
We added a short robustness experiment in Section 5.3: for one synthetic test (Branin-Currin) we ran qPOTS with and without additive Gaussian observation noise (for a range of variance values). We report hypervolume convergence curves (mean $\pm$ std over 10 repeats) showing that qPOTS maintains stable convergence relative to baselines up to the tested noise levels.
Figure 12: qPOTS on Branin–Currin under varying observation noise ($\sigma^{2}$). Curves show mean hypervolume across repetitions; shaded regions are $\pm 1$ standard deviation.